# auxmulticols[1]

### v3

@@ 1. Introduction
Machine Learning (ML), particularly deep learning, has gained widespread adoption in various real-life applications, such as face recognition and robotics, owing to its impressive performance. However, ML algorithms are susceptible to different types of attacks, such as model inversion (Fredrikson et al., 2015; Zhao et al., 2021) and membership inference (Shokri et al., 2017; Salem et al., 2019), which can jeopardize user privacy and lead to unintended information leakage. To overcome these challenges, federated learning (FL) (McMahan et al., 2017) has emerged as a promising solution, which enables collaborative model training between a central server and multiple clients while ensuring user-level privacy and reducing data fragmentation. Despite the advantages of FL, recent research has identified a new threat to its security: GAN-based feature inference attacks (Hitaj et al., 2017; Wang et al., 2019). In contrast to traditional feature inference attacks that exploit record-level privacy leakage to infer individual training images from trained models (Melis et al., 2019; Luo et al., 2021; Zhang et al., 2020a; Luo et al., 2022a), GAN-based attacks aim to infer the group-level information of private training images, i.e, the data distribution, which poses a significant challenge to existing privacy-preserving mechanisms such as differential privacy (DP) (Xiao et al., 2010). Although DP has been shown to be effective in preventing record-level privacy leakage in FL (Shokri et al., 2017), protecting group-level information is not guaranteed by its definition (Jiang et al., 2024b). This highlights the need for developing novel privacy-preserving techniques that can safeguard group-level privacy in FL. However, despite its significance, the current research landscape has yet to adequately explore this issue. It is worth noting that the existing privacy-preserving mechanisms (Scheliga et al., 2022; Gao et al., 2021) in FL are primarily designed to counteract record-level attacks (Zhu et al., 2019; Luo et al., 2021), and are insufficient in mitigating GAN-based attacks. Therefore, in this paper, we aim to address the vulnerability of FL to GAN-based attacks by developing a robust defense that is capable of mitigating the privacy risks posed by such attacks while effectively preserving the model utility.
@@
There are three primary challenges to consider when designing defenses against GAN-based attacks in the context of FL. First, traditional cryptographic methods, such as homomorphic encryption (Wu et al., 2020; Keller, 2020) and secure multiparty computation (Bonawitz et al., 2017; Mohassel and Zhang, 2017), offer limited utility. The reason is that these methods require significant computational resources (Wu et al., 2020; Zhang et al., 2020b; Truex et al., 2019), thereby restricting their application scenarios. In addition, encrypting local gradients can inhibit anomaly detection methods designed to counteract poisoning-based attacks (Jiang et al., 2024b; Bagdasaryan et al., 2020). Second, the effectiveness of noise injection, particularly techniques like DP (Xiao et al., 2010; Truex et al., 2019; Zhuang et al., 2024), is limited in mitigating GAN-based attacks, as demonstrated in previous research (Hitaj et al., 2017). While DP provides theoretical privacy guarantees for individual private images, GANs aim to reconstruct representations of specific classes (Triastcyn and Faltings, 2020; Xiao et al., 2018), surpassing the scope of DP’s privacy guarantee. Third, the success of GAN-based attacks is closely tied to the practicality of the federated model and the memorization nature of neural network models (Carlini et al., 2019). In other words, defending against such attacks becomes challenging as long as the victim’s local model maintains a sufficiently high classification accuracy (Hitaj et al., 2017).
@@
It is worth noting that existing defense mechanisms, such as cryptographic methods and noise injection, primarily focus on safeguarding privacy during the training phase, but their effectiveness in mitigating GAN-based attacks is limited. Motivated by this observation, we shift our attention towards privacy protection in the data generation phase before model training and propose a defense mechanism called Anti-GAN to address these above challenges. The core idea underlying Anti-GAN is to generate training images using a GAN model, where the visual characteristics of the generated images are obfuscated through a carefully designed loss function. This approach ensures that the attacker’s GAN model can only learn the distribution of visually indistinguishable images. However, as the distribution of the obfuscated images may differ from that of real images, this strategy could potentially result in a degradation of accuracy when the federated model is tested on real images. To overcome this issue, we devise a novel structure for the victim’s GAN model, enabling effective preservation of the classification features from real images while reasonably corrupting the visual features of the generated images. To further enhance the performance of the federated model, we devise a modified Mixup (Zhang et al., 2018) method to generate the final training images by mixing the generated fake images with the real images. It is important to note that Anti-GAN is specifically designed to obfuscate object details rather than the entire visual features of training images. Anti-GAN can preserve the semantic information necessary for the downstream classifier, such as the outline of a human face, while rendering object details, such as eyes and noses, indistinguishable to the human eye. Consequently, our scheme can significantly reduce the accuracy degradation caused to the federated model and ensure that images reconstructed by attackers are rendered useless.
@@
The framework of Anti-GAN is depicted in Fig. 1. Particularly, the victim initiates the defense process by training a GAN model using the private dataset $X$, generating corresponding fake images $X^{\prime}$. The objective is to ensure that $X^{\prime}$ and $X$ possess similar classification features while obscuring the visual features of $X^{\prime}$. Subsequently, the generated $X^{\prime}$ is mixed with $X$ to form the training dataset $\hat{X}$, which is then used to train the federated model. To assess the effectiveness of Anti-GAN, we conducted extensive experiments utilizing the MNIST (Yann et al., 1998), FashionMNIST (Xiao et al., 2017), and CelebA (Liu et al., 2015) datasets. The results demonstrate that Anti-GAN successfully reduces the plausibility of images reconstructed by the attacker’s GAN model while having minimal impact on the accuracy of the global federated model. Our results also show that, beyond countering group-level GAN-based attacks, Anti-GAN can effectively mitigate record-level gradient-matching attacks (Zhu et al., 2019) as well. Our contributions are summarized as follows:
@@
  • We conduct extensive experiments on three real-world datasets to comprehensively evaluate the performance of Anti-GAN. The results validate the efficacy of our scheme in protecting private images against GAN-based inference attacks.
@@ 5.6. Defense Performance under Different Client Settings
In this section, we evaluate the performance of Anti-GAN under different client configurations within the FL framework. The numbers of clients are chosen from $\{2,4,8,26,32\}$. Note that GAN-based attacks (Hitaj et al., 2017) are effective only in scenarios where the adversary does not possess the target data distribution, which means that the data distribution among different clients in FL should be heterogeneous, non-iid distributions. In the experiment, we first randomly divide the training data into two halves based on class distribution and similarly partition the clients into two groups. Then, each client in the first group is assigned a specific number of data samples randomly chosen from the first half of the training set, while clients in the second group receive samples from the second half. This approach ensures that the data distributions between the two client groups differ significantly, thereby enabling the attacker in one group to infer the data distribution of a defender in the opposite group using GAN-based attacks. To more accurately reflect real-world FL scenarios, we progressively reduce the number of samples per client as the number of clients increases. Specifically, the sample sizes per client for client configurations $\{2,4,8,26,32\}$ are set to $\{6000,4000,2000,1000,500\}$, respectively.
@@ 7. Related Work
Inference Attacks on Machine Learning. Machine Learning (ML) algorithms are vulnerable to various forms of inference attacks, including membership inference (Shokri et al., 2017; Salem et al., 2019; Nasr et al., 2019; Melis et al., 2019), property inference (Wang et al., 2022; Ganju et al., 2018; Mahloujifar et al., 2022; Melis et al., 2019; Jiang et al., 2024b), and feature inference (also known as model inversion) (Fredrikson et al., 2015; Luo et al., 2021; Wu et al., 2016; Zhu et al., 2019; Hitaj et al., 2017; Wang et al., 2019; Luo et al., 2022a; Jayaraman and Evans, 2022). Membership inference attacks determine whether a specific sample is part of the training dataset. (Shokri et al., 2017) introduces a shadow model training technique for membership inference, which is later improved by (Salem et al., 2019) under loose assumptions, and by (Nasr et al., 2019) in the setting of federated learning. Property inference attacks aim to infer aggregate statistics of the training dataset, such as the statistics of software execution traces (Ganju et al., 2018) and the age distribution of a bank’s customer data (Jiang et al., 2024b). Feature inference attacks attempt to reconstruct the original inputs from ML applications. (Fredrikson et al., 2015) explores model inversion attacks in decision trees and neural networks using labels and black-box access to the target model. (Luo et al., 2021) proposes to infer the training records in federated learning via model predictions. (Zhu et al., 2019) restores the private training data using only model gradients. In federated learning, recent studies exploit feature inference attacks based on GANs. (Hitaj et al., 2017) proposes a GAN-based threat model where an adversary tries to extract representatives of a class not owned by him. By taking the shared global model as the discriminator of GAN, the adversary can further mislabel the pseudo-samples generated by the generator to trick other participants into leaking more private information. (Wang et al., 2019) assumes a malicious server and utilizes GANs to explore user-level privacy leakage, similar to the method in (Hitaj et al., 2017). Note that in the literature, GANs have also been used in other types of attacks, such as membership inference attacks (Chen et al., 2020; Lin et al., 2021), backdoor attacks (Mei et al., 2023; Hu and Tan, 2022), and adversarial attacks (Zhu et al., 2023; Na et al., 2022). In this paper, we focus on GAN-based feature inference attacks. The main difference between GAN-based feature inference attacks (Hitaj et al., 2017; Wang et al., 2019) and traditional feature inference attacks (Fredrikson et al., 2015; Luo et al., 2021; Wu et al., 2016; Zhu et al., 2019; Jayaraman and Evans, 2022) is that the former focuses on inferring class representations (i.e., group-level information) via a GAN model from model gradients, whereas the latter aims to restore record-level private information (i.e., individual model inputs) from publicly released information, such as data labels (Fredrikson et al., 2015), model predictions (Luo et al., 2021), and gradients (Zhu et al., 2019). Record-level feature inference attacks can be effectively mitigated by state-of-the-art defense mechanisms in FL, such as dropout (Melis et al., 2019; Salem et al., 2019), differential privacy (Shokri and Shmatikov, 2015; Bagdasaryan et al., 2020), and secure multiparty computation (Bonawitz et al., 2017). However, GAN-based feature inference attacks pose significant challenges for mitigation. The effectiveness of these attacks is closely tied to the practicality of the federated model (Hitaj et al., 2017), making them difficult to defend against. To the best of our knowledge, there are no defense mechanisms that can effectively counter GAN-based attacks without causing substantial performance degradation to the federated model. In this paper, we focus on developing defenses against GAN-based feature inference attacks (Hitaj et al., 2017; Wang et al., 2019) that can achieve a favorable privacy-utility trade-off.
@@
Defenses against Inference Attacks. Numerous defense methods have been proposed to mitigate the risks incurred by inference attacks, including dropout (Melis et al., 2019; Salem et al., 2019; Scheliga et al., 2023), model stacking (Salem et al., 2019), mutual information-based regularization (Wang et al., 2021), anomaly detection (Bagdasaryan et al., 2020), differential privacy (Shokri and Shmatikov, 2015; Bagdasaryan et al., 2020; Geyer et al., 2017; Jiang et al., 2024a; Guo et al., 2022), secure multiparty computation (Wu et al., 2020; Mohassel and Zhang, 2017; Bonawitz et al., 2017; Zhang et al., 2020b; Truex et al., 2019), and InstaHide (Huang et al., 2020). Dropout (Melis et al., 2019; Salem et al., 2019; Scheliga et al., 2023) and model stacking (Salem et al., 2019) are basically regularization methods, which are mainly used to prevent model overfitting, thereby reducing the amount of information memorized by ML models. Anomaly detection (Bagdasaryan et al., 2020) is utilized to mitigate poisoning-based attacks performed by active adversaries. Traditional differential privacy methods (Shokri and Shmatikov, 2015; Bagdasaryan et al., 2020; Guo et al., 2022) provide rigorous theoretical guarantees for privacy protection, but they are applicable to individual-level attacks instead of group-level attacks. Although recent studies (Geyer et al., 2017; Jiang et al., 2024a) have extended DP guarantees to group-level privacy, these methods can significantly degrade model performance. Secure multiparty computation (Wu et al., 2020; Mohassel and Zhang, 2017; Bonawitz et al., 2017) can impose substantial computational costs during model training and has not been widely applied in real-world applications. Among these defense methods, InstaHide (Huang et al., 2020) is most similar to our approach, i.e., using Mixup (Zhang et al., 2018) to corrupt the visual features of private training images. However, there are three main differences between InstaHide and Anti-GAN. First, InstaHide mixes multiple real images into one training image, while our scheme uses a GAN model to generate multiple fake images and then mixes these fake images with real images. Second, InstaHide obfuscates image visual features by randomly flipping pixel signs, leading to uncontrollable model performance degradation. In contrast, Anti-GAN obfuscates visual features using a specifically designed loss function, allowing the privacy-utility tradeoff to be tuned for different application scenarios. Third, (Carlini et al., 2021) and (Luo et al., 2022b) have experimentally demonstrated that private images can be reconstructed from the mixup images generated by InstaHide on the condition that the same private image is mixed into multiple training images. Our scheme mitigates this risk by ensuring that each private image is mixed only once. It is important to note that none of the aforementioned defenses, including InstaHide, are specifically designed to counter GAN-based attacks. To the best of our knowledge, this is the first paper to exploit defenses against GAN-based feature inference attacks in the context of federated learning.

### v4

@@ 1. Introduction
Machine Learning (ML), particularly deep learning, has gained widespread adoption in various real-life applications, such as face recognition and robotics, owing to its impressive performance. However, ML algorithms are susceptible to different types of attacks, such as model inversion (Fredrikson et al., 2015; Zhao et al., 2021) and membership inference (Shokri et al., 2017; Salem et al., 2019), which can jeopardize user privacy and lead to unintended information leakage. To overcome these challenges, federated learning (FL) (McMahan et al., 2017) has emerged as a promising solution, which enables collaborative model training between a central server and multiple clients while ensuring user-level privacy and reducing data fragmentation. Despite the advantages of FL, recent research has identified a new threat to its security: GAN-based feature inference attacks (Hitaj et al., 2017; Wang et al., 2019). In contrast to traditional feature inference attacks that exploit record-level privacy leakage to infer individual training images from trained models (Melis et al., 2019; Luo et al., 2021; Zhang et al., 2020a; Luo et al., 2022a), GAN-based attacks aim to infer the group-level information of private training images, i.e, the data distribution. This presents a critical challenge to existing privacy-preserving mechanisms (Scheliga et al., 2022; Gao et al., 2021) in FL, as they are primarily designed to counteract record-level attacks (Zhu et al., 2019; Luo et al., 2021) and are often insufficient for defending against GAN-based attacks. Despite its significance, the current research landscape has yet to adequately explore this issue. In this paper, we aim to address the vulnerability of FL to GAN-based attacks by developing a robust defense that is capable of mitigating the privacy risks posed by such attacks while effectively preserving the model utility.
@@
There are three primary challenges to consider when designing defenses against GAN-based attacks in the context of FL. First, traditional cryptographic methods, such as homomorphic encryption (Wu et al., 2020; Keller, 2020) and secure multiparty computation (Bonawitz et al., 2017; Mohassel and Zhang, 2017), offer limited utility. The reason is that these methods require significant computational resources (Wu et al., 2020; Zhang et al., 2020b; Truex et al., 2019), thereby restricting their application scenarios. In addition, encrypting local gradients can inhibit anomaly detection methods designed to counteract poisoning-based attacks (Jiang et al., 2024b; Bagdasaryan et al., 2020). Second, the effectiveness of noise injection, particularly techniques like DP (Xiao et al., 2010; Truex et al., 2019; Zhuang et al., 2024), is limited in mitigating GAN-based attacks (Hitaj et al., 2017; Jiang et al., 2024b). While DP provides theoretical privacy guarantees for individual private images, GANs aim to reconstruct representations of specific classes (Triastcyn and Faltings, 2020; Xiao et al., 2018), surpassing the initial scope of DP’s privacy guarantee. Moreover, recent work (Jiang et al., 2024b) has demonstrated the ability to extract training data distributions from FL systems protected by DP. While extensions of DP to group-level privacy have been explored (Geyer et al., 2017; Jiang et al., 2024a), these approaches often compromise model performance under moderate privacy budgets (Geyer et al., 2017), and the application scenarios are typically limited to simple models like logistic regression or support vector machines (Jiang et al., 2024a). Third, the success of GAN-based attacks is closely tied to the practicality of the federated model and the memorization nature of neural network models (Carlini et al., 2019). In other words, defending against such attacks becomes challenging as long as the victim’s local model maintains a high classification accuracy (Hitaj et al., 2017).
@@
It is worth noting that existing defense mechanisms, such as cryptographic methods and noise injection, primarily focus on safeguarding privacy during the training phase. Considering their limited effectiveness in mitigating GAN-based attacks, we shift our attention for privacy protection from the model training phase to the data generation phase. Correspondingly, we propose a defense mechanism called Anti-GAN to protect the group-level information against GAN-based attacks. The core idea underlying Anti-GAN is to generate training images using a GAN model, where the visual characteristics of the generated images are obfuscated through a carefully designed loss function. This approach ensures that the attacker’s GAN model can only learn the distribution of visually indistinguishable images. However, as the distribution of the obfuscated images may differ from that of real images, this strategy could potentially result in a degradation of accuracy when the federated model is tested on real images. To overcome this issue, we devise a novel structure for the victim’s GAN model, enabling effective preservation of the classification features from real images while reasonably corrupting the visual features of the generated images. To further enhance the performance of the federated model, we devise a modified Mixup (Zhang et al., 2018) method to generate the final training images by mixing the generated fake images with the real images. It is important to note that Anti-GAN is specifically designed to obfuscate object details rather than the entire visual features of training images. Anti-GAN can preserve the semantic information necessary for the downstream classifier, such as the outline of a human face, while rendering object details, such as eyes and noses, indistinguishable to the human eye. Consequently, our scheme can significantly reduce the accuracy degradation caused to the federated model and ensure that images reconstructed by attackers are rendered useless.
@@
The framework of Anti-GAN is depicted in Fig. 1. Particularly, the victim initiates the defense process by training a GAN model using the private dataset $X$, generating corresponding fake images $X^{\prime}$. The objective is to ensure that $X^{\prime}$ and $X$ possess similar classification features while obscuring the visual features of $X^{\prime}$. Subsequently, the generated $X^{\prime}$ is mixed with $X$ to form the training dataset $\hat{X}$, which is then used to train the federated model. To assess the effectiveness of Anti-GAN, we conducted extensive experiments utilizing the MNIST (Yann et al., 1998), FashionMNIST (Xiao et al., 2017), CelebA (Liu et al., 2015), and CIFAR10 (Krizhevsky et al., 2009) datasets. The results demonstrate that Anti-GAN successfully reduces the plausibility of images reconstructed by the attacker’s GAN model while having minimal impact on the accuracy of the global federated model. Our results also show that, beyond countering group-level GAN-based attacks, Anti-GAN can effectively mitigate record-level gradient-matching attacks (Zhu et al., 2019) as well. Our contributions are summarized as follows:
@@
  • We conduct extensive experiments on four real-world datasets to comprehensively evaluate the performance of Anti-GAN. The results validate the efficacy of our scheme in protecting private images against GAN-based inference attacks.
@@ 5.6. Defense Performance under Different Client Settings
In this section, we evaluate the performance of Anti-GAN under different client configurations within the FL framework. The numbers of clients are chosen from $\{2,4,8,{\color[rgb]{0,0,0}16},32\}$. Note that GAN-based attacks (Hitaj et al., 2017) are effective only in scenarios where the adversary does not possess the target data distribution, which means that the data distribution among different clients in FL should be heterogeneous, non-iid distributions. In the experiment, we first randomly divide the training data into two halves based on class distribution and similarly partition the clients into two groups. Then, each client in the first group is assigned a specific number of data samples randomly chosen from the first half of the training set, while clients in the second group receive samples from the second half. This approach ensures that the data distributions between the two client groups differ significantly, thereby enabling the attacker in one group to infer the data distribution of a defender in the opposite group using GAN-based attacks. To more accurately reflect real-world FL scenarios, we progressively reduce the number of samples per client as the number of clients increases. Specifically, the sample sizes per client for client configurations $\{2,4,8,{\color[rgb]{0,0,0}16},32\}$ are set to $\{6000,4000,2000,1000,500\}$, respectively.
@@ 7.1. Inference Attacks on Machine Learning
Machine Learning (ML) algorithms are vulnerable to various forms of inference attacks, including membership inference (Shokri et al., 2017; Salem et al., 2019; Nasr et al., 2019; Melis et al., 2019), property inference (Wang et al., 2022; Ganju et al., 2018; Mahloujifar et al., 2022; Melis et al., 2019; Jiang et al., 2024b; Luo et al., 2024), and feature inference (also known as model inversion) (Fredrikson et al., 2015; Luo et al., 2021; Wu et al., 2016; Zhu et al., 2023a, 2019; Hitaj et al., 2017; Wang et al., 2019; Luo et al., 2022a; Jayaraman and Evans, 2022). Membership inference attacks determine whether a specific sample is part of the training dataset. (Shokri et al., 2017) introduces a shadow model training technique for membership inference, which is later improved by (Salem et al., 2019) under loose assumptions, and by (Nasr et al., 2019) in the setting of federated learning. Property inference attacks aim to infer aggregate statistics of the training dataset, such as the statistics of software execution traces (Ganju et al., 2018) and the age distribution of a bank’s customer data (Jiang et al., 2024b). Feature inference attacks attempt to reconstruct the original inputs from ML applications. (Fredrikson et al., 2015) explores model inversion attacks in decision trees and neural networks using labels and black-box access to the target model. (Luo et al., 2021) proposes to infer the training records in federated learning via model predictions. (Zhu et al., 2019) restores the private training data using only model gradients. In federated learning, recent studies exploit GAN-based feature inference attacks. (Hitaj et al., 2017) proposes a GAN-based threat model where an adversary tries to extract representatives of a class not owned by him. By taking the shared global model as the discriminator of GAN, the adversary can further mislabel the pseudo-samples generated by the generator to trick other participants into leaking more private information. (Wang et al., 2019) assumes a malicious server and utilizes GANs to explore user-level privacy leakage, similar to the method in (Hitaj et al., 2017). Note that in the literature, GANs have also been used in other types of attacks, such as membership inference attacks (Chen et al., 2020; Lin et al., 2021), backdoor attacks (Mei et al., 2023; Hu and Tan, 2022), and adversarial attacks (Zhu et al., 2023b; Na et al., 2022).
In this paper, we focus on GAN-based feature inference attacks. The main difference between GAN-based feature inference attacks (Hitaj et al., 2017; Wang et al., 2019) and traditional feature inference attacks (Fredrikson et al., 2015; Luo et al., 2021; Wu et al., 2016; Zhu et al., 2019; Jayaraman and Evans, 2022) is that the former focuses on inferring class representations (i.e., group-level information) via a GAN model from model gradients, whereas the latter aims to restore record-level private information (i.e., individual model inputs) from publicly released information, such as data labels (Fredrikson et al., 2015), model predictions (Luo et al., 2021), and gradients (Zhu et al., 2019). Record-level feature inference attacks can be effectively mitigated by state-of-the-art defense mechanisms in FL, such as dropout (Melis et al., 2019; Salem et al., 2019), differential privacy (Shokri and Shmatikov, 2015; Bagdasaryan et al., 2020), and secure multiparty computation (Bonawitz et al., 2017). However, GAN-based feature inference attacks pose significant challenges for mitigation. The effectiveness of these attacks is closely tied to the practicality of the federated model (Hitaj et al., 2017), making them difficult to defend against. To the best of our knowledge, there are no defense mechanisms that can effectively counter GAN-based attacks without causing substantial performance degradation to the federated model. In this paper, we focus on developing defenses against GAN-based feature inference attacks (Hitaj et al., 2017; Wang et al., 2019) that can achieve a favorable privacy-utility trade-off.
7.2. Defenses against Inference Attacks
@@
Numerous defense methods have been proposed to mitigate the risks incurred by inference attacks, including dropout (Melis et al., 2019; Salem et al., 2019; Scheliga et al., 2023), model stacking (Salem et al., 2019), anomaly detection (Bagdasaryan et al., 2020), differential privacy (Shokri and Shmatikov, 2015; Bagdasaryan et al., 2020; Geyer et al., 2017; Jiang et al., 2024a; Guo et al., 2022), secure multiparty computation (Wu et al., 2020; Mohassel and Zhang, 2017; Bonawitz et al., 2017; Zhang et al., 2020b; Truex et al., 2019), and InstaHide (Huang et al., 2020). Dropout (Melis et al., 2019; Salem et al., 2019; Scheliga et al., 2023) and model stacking (Salem et al., 2019) are basically regularization methods, which are mainly used to prevent model overfitting, thereby reducing the amount of information memorized by ML models. Anomaly detection (Bagdasaryan et al., 2020) is utilized to mitigate poisoning-based attacks performed by active adversaries. Traditional differential privacy methods (Shokri and Shmatikov, 2015; Bagdasaryan et al., 2020; Guo et al., 2022) provide rigorous theoretical guarantees for privacy protection, but they are applicable to individual-level attacks instead of group-level attacks. Although recent studies (Geyer et al., 2017; Jiang et al., 2024a, c) have extended DP guarantees to group-level privacy, these methods can significantly degrade model performance. Secure multiparty computation (Wu et al., 2020; Mohassel and Zhang, 2017; Bonawitz et al., 2017) can impose substantial computational costs during model training and has not been widely applied in real-world applications.
Among these defense methods, InstaHide (Huang et al., 2020) is most similar to our approach, i.e., using Mixup (Zhang et al., 2018) to corrupt the visual features of private training images. However, there are three main differences between InstaHide and Anti-GAN. First, InstaHide mixes multiple real images into one training image, while our scheme uses a GAN model to generate multiple fake images and then mixes these fake images with real images. Second, InstaHide obfuscates image visual features by randomly flipping pixel signs, leading to uncontrollable model performance degradation. In contrast, Anti-GAN obfuscates visual features using a specifically designed loss function, allowing the privacy-utility tradeoff to be tuned for different application scenarios. Third, (Carlini et al., 2021) and (Luo et al., 2022b) have experimentally demonstrated that private images can be reconstructed from the mixup images generated by InstaHide on the condition that the same private image is mixed into multiple training images. Our scheme mitigates this risk by ensuring that each private image is mixed only once. It is important to note that none of the aforementioned defenses, including InstaHide, are specifically designed to counter GAN-based attacks. To the best of our knowledge, this is the first paper to exploit defenses against GAN-based feature inference attacks in the context of federated learning.